OpenStudy (anonymous):

solve for y 2x+15y=19

jigglypuff314 (jigglypuff314):

subtract 2x from both sides divide both sides by 15

OpenStudy (anonymous):

so 14 divided by 17?

OpenStudy (anonymous):

i mean 13*

jigglypuff314 (jigglypuff314):

2x+15y=19 subtract 2x from both sides -2x -2x 15y = -2x + 19 divide both sides by 15 /15 /15 /15 y = -2/15 x + 19/15

OpenStudy (anonymous):

thnks :)

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

can you help me with another one?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Solve the following proportion for v.  round your answer to the nearest tenth

jigglypuff314 (jigglypuff314):

v = 8 * 13 / 11

OpenStudy (anonymous):

thanks so much!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

I mean 9.454545555

jigglypuff314 (jigglypuff314):

to the nearest tenth though right? 4.4545 ^tenths place so 4.5

OpenStudy (anonymous):

thanks again :)